OpenStudy (anonymous):

What is the solution to the equation 21 x − 3 = 14?

OpenStudy (compassionate):

Hi, You want to isolate x. So you have 21x - 3 = 14 We can start by adding 3 to both sides, which gives us: 21x = 17 Then you would simply divide by 21.

OpenStudy (anonymous):

the x-3 is supposed to be the power of 21

OpenStudy (compassionate):

Could you type your problem using the "Equation" button located at the bottom left?

OpenStudy (anonymous):

ok ill do that

OpenStudy (anonymous):

What is the solution to the equation 21^ x − 3 = 14?

OpenStudy (compassionate):

\[21^{x-3} = 14\] Is that it?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Thank you @compassionate